OpenStudy (anonymous):

Assuming that all matrices are n x n and invertible, solve for D ?? A B (A^T) D B (C^T) C = A B^T.

OpenStudy (kinggeorge):

I'm assuming \(A^T\) is the transpose of \(A\)?

OpenStudy (kinggeorge):

If so, just write out what you have. \[A \cdot B \cdot A^T \cdot D \cdot B \cdot C^T \cdot C=A B^T\]

OpenStudy (kinggeorge):

Then start multiplying by inverses on either side until you only have \(D\) on the left side. \[(A B A^T)^{-1} \cdot (ABA^T) \cdot D \cdot (BC^TC)(BC^TC)^{-1}=D=\]\[(ABA^T)^{-1} AB^T (BC^TC)^{-1}\]

OpenStudy (kinggeorge):

Did this make sense?

OpenStudy (anonymous):

\[ D = B^{-1} A^{-T} B^{T} B^{-1} C^{-T} C^{-1} This is true ?? =(\]

OpenStudy (kinggeorge):

Close. Rather, \[D=(A^T)^{-1} B^{-1} B^T C^{-1} (C^T)^{-1}B^{-1}\]Remember, if you have three invertible matrices \(X, Y, Z\) then\[(XYZ)^{-1} = Z^{-1}Y^{-1}X^{-1}\]

OpenStudy (anonymous):

\[OK ,, D = ( B A^{T} B C^{T} C )^{-1} B^{T} \]

OpenStudy (kinggeorge):

No... \[(BA^TBC^TC)^{−1}=C^{-1} (C^T)^{-1} B^{-1}(A^T)^{-1} B^{-1}\]This is not the same result as what we obtained above.

OpenStudy (anonymous):

why? ="" this is same =$

OpenStudy (phi):

King has it correct \[ D=(A^T)^{-1} B^{-1} B^T C^{-1} (C^T)^{-1}B^{-1} \] The easiest way to get this is to premultiply on the left: multiply both sides on the left by A^-1, then B^-1 then A^-T (inverse and transpose) post multiply on the right in the order C^-1 the C^-T then B^-1

OpenStudy (anonymous):

Thx king and phi I understand what you are saying but the issue from the outset so A B (A^T) D B (C^T) C = A B^T. First : (A A^-1) B (A^T) D B (C^T) C = (AA^-1) B^T : B (A^{T}) D B (C^{T}) C = B^{T} : (BB^{-1}) (A^{T} A^{-T}) D B C^{T} C = B^{T} B^{-1} A^{-T} : D B (C^{T}) C = B^{T} B^{-1} A^{-T} : D (C^{T} C^{-T}) (C C^{-1}) (B B^{-1}) = B^{T} B^{-1} A^{-T} C^{-T} C^{-1} B^{-1} : D = B^{T} B^{-1} A^{-T} C^{-T} C^{-1} B^{-1} : D = ( B C C^{T} A^{T} B )^{-1} B^{T} This is reached =$

OpenStudy (phi):

As you know matrix multiplication is (in general) not commutative. That is AB≠BA So you must be careful when multiplying, to always multiply on the same side: So it is more correct to write ( A^-1A) B (A^T) D B (C^T) C = (A^-1A) B^T (A^-1 goes on the left) than what you wrote:(A A^-1) B (A^T) D B (C^T) C = (AA^-1) B^T Going from : B (A^{T}) D B (C^{T}) C = B^{T} to : (BB^{-1}) (A^{T} A^{-T}) D B C^{T} C = B^{T} B^{-1} A^{-T} is WRONG. (You cannot switch the order of B^T and B^-1 on the right-hand side.)

OpenStudy (anonymous):

أهاا ,,لم انتبه لذلك اشكرك :(

OpenStudy (anonymous):

A B (A^T) D B (C^T) C = A B^T. (A^{-1} A) B (A^{T}) D B (C^{T}) C = (A^{-1}A) B^{T} (B^{-1} B) A^{T} D B (C^{T}) C = B^{T} B^{-1} (A^-1)^{T} A^{T} D B (C^{T}) C = (A^-1)^{T} B^{T} B^{-1} D (B^{-1}B) (C^{T}) C = (A^-1)^{T} B^{T} B^{-1} B^{-1} D (C^-1)^{T} (C^{T}) C = (A^-1)^{T} B^{T} B^{-1} B^{-1} (C^T)^{-1} D C^{-1} C = (A^-1)^{T} B^{T} B^{-1} B^{-1} (C^T)^{-1} C^{-1} D = (A^-1)^{T} B^{T} B^{-1} B^{-1} (C^T)^{-1} C^{-1} <<I hope that to be true phi

OpenStudy (phi):

On your 2nd line you start with B (A^{T}) D B (C^{T}) C = B^{T} (where I have removed A^-1 A because that is just I, the identity matrix) you then wrote (B^{-1} B) A^{T} D B (C^{T}) C = B^{T} B^{-1} BUT this is not correct. You multiply B^-1 on the left side (you pre-multiplied) on the LHS (left hand side of the equal sign). But on the RHS of the equal you post-multiplied by B^-1. You should have written (B^{-1} B) A^{T} D B (C^{T}) C = B^{-1}B^{T} Remember order matters. If you start with the equation A=A and then multiply both sides by B: BA= BA is correct, but BA=AB is not correct.